OpenStudy (anonymous):

Show that 1+kx is the local linearization of (1+x)^k near x = 0

OpenStudy (anonymous):

take the derivative, replace x by 0 to find your slope, then use point - slope formula

OpenStudy (anonymous):

\[f(x)=(1+x)^k\] \[f'(x)=k(1+x)^{k-1}\] \[f'(0)=k\] \[y-y_1=m(x=x_1)\] \[y-1=k(x-0)\] \[y=1+kx\]